OpenStudy (dls):

A wire of length L and mass M is bent in the form of a rectangle ABCD with AB/BC=2.The Moment of inertia of the wireframe about the side BC is?

OpenStudy (yrelhan4):

solve krke batata hun.

OpenStudy (yrelhan4):

haha abhi hua nhi. thoda confuse ho gya hun. :/

OpenStudy (shubhamsrg):

^_^

OpenStudy (dls):

<3

OpenStudy (yrelhan4):

hmm. 2*m/3(l/6)^2 + m/6(l/3)^2 .. solve kr le. aasan to tha :/

OpenStudy (shubhamsrg):

forget him parth, he is mine! :|

OpenStudy (dls):

\[\LARGE \frac{2m \times 36 }{3L^2}+\frac{9 \times m}{6L^{2}}\] ??

OpenStudy (yrelhan4):

ravi raj kon hai? :O ye kya kr rha hai? :O 2ml^2/(36*3) + ml^2(6*9)

OpenStudy (dls):

\[\LARGE \frac{3ML^2}{54}????\]

OpenStudy (dls):

@yrelhan4 ?:O

OpenStudy (yrelhan4):

ml^2/27 ??

OpenStudy (dls):

Kaafi easy tha khair.

OpenStudy (yrelhan4):

thik hai?

OpenStudy (dls):

No :P

OpenStudy (dls):

lol

OpenStudy (yrelhan4):

oh teri.... kya galti hai?

OpenStudy (dls):

IDK,answer is weird

OpenStudy (dls):

\[\LARGE \frac{11}{252}ML^2||\frac{8}{203}ML^2||\frac{5}{36}ML^2||\frac{7}{162}ML^2\] options^^

sam (.sam.):

 \[I_{bc}=\frac{mh^2}{3}+\frac{mw^2}{12} \\ \\ I_{bc}=m(\frac{h^2}{3}+\frac{4h^4}{12}) \\ \\ I_{bc}=m(\frac{h^2}{3}+\frac{h^4}{4})\] \[L=2w+2h \\ \\ h=\frac{L-2w}{2}\] hmm

OpenStudy (dls):

lol

sam (.sam.):

\[I_{bc}=\frac{mh^2}{3}+\frac{mw^2}{12} \\ \\ I_{bc}=m(\frac{h^2}{3}+\frac{4h^2}{12}) \\ \\ I_{bc}=m(\frac{h^2}{3}+\frac{h^2}{3}) \\ \\ I_{bc}=m(\frac{2h^2}{3})\] \[L=2w+2h \\ \\ h=\frac{L-2w}{2} \\ \\ \text{from}~ w=2h \\ \\ h=\frac{L}{6}\] \[I=\left( \left(\frac{2}{3} \left(\frac{L}{6}\right)^2\right)\right) m\] \[I=\frac{mL^2 }{54}\]

OpenStudy (dls):

Not in options..:(

OpenStudy (shubhamsrg):

 I(AB) about center of BC = (m/3)(l/3)^2 /12 + (m/3)(l/6)^2 + (m/3)(l/12)^2 I(DC) about center of BC = same as of AB I(AD) about center of BC = (m/6)(l/6^2) /12 + (m/6)(l/3)^2 I(BC) about center = (m/6)(l/6)^2 /12 I net about BC perp. to the plane = sum of all = 7/144 ml^2 so about BC = 7/288 ml^2 hmm, where did I go wrong ?